# SAR-Net: Multi-scale Direction-aware SAR Network via Global Information Fusion

### v1

@@ I Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring[1, 2]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships. In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to designing object detection methods using Convolutional Neural Networks (CNNs), yielding notable results [3, 4, 5, 6, 7, 8, 9]. However, the unique challenges posed by SAR images, characterized by their multi-scale and rotational nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands at a critical challenge in this field. Recent advancements in SAR image object detection have prominently featured deep learning approaches. Addressing the challenges of detecting objects in varying orientations, Li et al. [10] developed the Spatial Frequency Feature Fusion Network. This innovative network utilizes a Feature Pyramid Network (FPN) [11] to create spatial multi-scale features, specifically for ship objects, through a top-down hierarchical approach. Additionally, it employs a polar coordinate Fourier transform to attain rotation-invariant features in the frequency domain. In another notable work, Zhang et al. [12] introduced the Quad-Feature Pyramid Network (Quad-FPN), an architectural innovation consisting of four feature pyramid networks. This design effectively confronts issues such as background noise, multi-scale object representation, and the subtle features of small boats. Complementing these advancements, Li et al. [13] developed the Adjacent Feature Fusion (AFF) module, which strategically integrates local shallow features into adjacent layers. This selective and adaptive approach significantly enhances feature fusion, leading to improved representation of objects in SAR images.
@@
Current research in SAR image object detection primarily focuses on utilizing a FPN or its variants for feature fusion, offering a partial solution to multi-scale challenges. However, a significant limitation of these methods is their lack of global information integration at each layer. As illustrated in Fig. 1 (right side), local information flow between adjacent layers leads to a dilution of information across non-adjacent layers, resulting in information loss [14]. Recognizing the importance of global information, Bai et al. [15] proposed a Global Context-Guided Feature Balance Pyramid (GC-FBP), incorporating a global context module to bridge distant relationships between ship objects and their background. Yet, their approach of simply stacking context information onto the original feature map impedes effective interplay of information. Our method, depicted on the left side of Fig. 1, diverges from traditional algorithms that primarily depend on straightforward concatenation of adjacent feature layers. Instead, we adopt a global information fusion strategy, allowing for feature integration across various scales. This approach not only enhances the effectiveness and efficiency of the fusion process but also minimally impacts information loss during transmission. Traditional SAR object detectors, which predominantly use horizontal boxes, typically capture features of directional objects through an attention module [16] or by transforming features into the frequency domain [10]. However, these methods tend to neglect the correlation between the object and its contextual environment. This oversight often leads to ambiguous differentiation between the object and the background, with the resultant features being significantly noisy. Furthermore, these techniques generally overlook the crucial role of positional information. In contrast, our approach consider not only the object’s features but also the surrounding context. By integrating directional and positional information and embedding channel attention, our approach offers a comprehensive solution. We have distilled the problem of SAR image object detection into a broader challenge: addressing the issues of multi-scale and directional objects in predominantly background-dominated images. We base our considerations on three key premises: (1) the significant variation in object scales in SAR images. (2) the need to consider both shallow textural and deep semantic information of objects. (3) the directional nature of objects coupled with background interference. To address these challenges, we introduce a unity compensation mechanism, which includes three components: the Multi-scale Alignment Module (MAM), the Multi-level Fusion Module (MFM), and the Multi-feature Embedding Module (MEM). MAM initially unites multi-scale features across different spatial hierarchical levels. MFM then globally merges high-level semantic information with low-level detailed information, providing a comprehensive overview of the united features. Lastly, MEM compensates for any original feature to enrich their representational information. This multifaceted approach not only improves the detection of multi-scale objects but also significantly enhances information fusion capabilities. In addressing the third premise, we present a direction-aware attention module equipped with bidirectional attention polymerization capabilities. This innovative module employs deformable convolution and global pooling operations to extract directional and positional information along both vertical and horizontal axes of the input features. It then generates a direction-aware attention map, which is effectively integrated with the original input features. This approach adeptly tackles the challenge of accurately discerning directional and positional information amidst intricate background interference. Our contributions can be summarized in three primary points: $\bullet$ We introduce SAR-Net, a novel framework specifically designed for SAR image object detection. SAR-Net incorporates a unity compensation mechanism that globally unites multi-scale features of objects while compensating for the original feature. This ensures that global information is inherently integrated within features of different scales. $\bullet$ We design a direction-aware attention module with bidirectional attention polymerization capabilities. This module extracts object features while obtaining directional and positional information, enabling the identification of object rotation poses. Such a design is particularly effective in complex SAR images, where the directional attributes of objects are often obscured by significant background interference. $\bullet$ To validate the effectiveness, generalizability, and robustness of SAR-Net, we conduct comprehensive comparative and ablation experiments using three publicly available benchmark datasets for SAR aircraft and ship detection. The results of these experiments clearly show that our SAR-Net framework outperforms other state-of-the-art methods in all three datasets. The remainder of this paper is organized as follows. Section II reviews the development of object detection in the field of SAR images and related literature on solving multi-scale problems. Section III details our proposed framework and modules. In Section IV, the experimental setup is described, and qualitative and quantitative results are reported. Finally, conclusions are drawn in Section V.
@@ II-A SAR Object Detection
Over the past few years, object detection in SAR images has aroused rising attention. Traditional detection algorithms, adhering to Constant False Alarm Rate (CFAR) principles [17], operate on a statistical model, dynamically setting the detection threshold based on a specified false alarm rate and the statistical distribution of background clutter. Consequently, a multitude of CFAR-based algorithms have surfaced. In a recent work [18], a novel decomposition approach was introduced to scrutinize the scattering between ships and the sea background. This decomposition technique, when integrated with CFAR, proves effective in ship detection. In a different vein, Lang et al. [19] enhanced the representation of pixels by incorporating spatial and intensity information, markedly improving the distinguishability between ships and background elements. However, the advent of high-resolution SAR images has introduced challenges such as complex ground interference and textured scenes accompanying multi-scale objects. This complexity renders the calculation of significant background clutter pixels in large-scale SAR images time-consuming. In response to the challenges of multi-scale object detection in SAR images, scholars have proposed several refined CFAR object detection algorithms. Examples include the scale sliding window object detection algorithm [20], the optimized super-pixel CFAR object detection algorithm [21], and the iterative CFAR object detection algorithm [22]. For detecting objects at larger scales, Zhai et al. [23] devised a ship detection algorithm emphasizing saliency and contextual information processing. This approach excels in focusing on sizable ships and background objects with prominent features, yielding commendable results. Despite these achievements, traditional methods still exhibit limitations in detecting small objects, necessitate manual parameter configuration, and heavily rely on manual design, presenting challenges for robustness and hindering their applicability to other SAR datasets. Consequently, there is a pressing need for more adaptive and automated approaches in SAR object detection. Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies. Shifting focus to SAR ship detection and identification, Zhang et al. [24] proposed a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of SAR aircraft detection and recognition, Zhao et al. [25] introduced an attention module, strategically fusing refined low-level texture features with high-level semantic features. This fusion strategy enhances aircraft detection rates by leveraging complementary information. Furthermore, Zhao et al. [26] devised a multi-branch dilated convolution feature pyramid method. This innovative approach minimizes redundancy while accentuating critical aircraft features through the establishment of dense connections. In a related endeavor, Zhao et al. [27] put forth a multitask learning framework for object detection (MLDet) tailored for ship detection in SAR images. This end-to-end framework encompasses tasks such as object detection, speckle suppression, and object segmentation. The collaborative execution of these tasks contributes to an enhanced detection performance. Addressing the specific challenge of large-scale SAR images, Yang et al. [28] concentrated on a multi-channel aggregation framework. This framework holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition. Although progress has been made, scale variation across object instances is still one of the most knotty problems in SAR object detection.
@@ II-B Multi-scale Object Detection
Within the domain of deep learning, the application of image pyramids initially emerged as a solution to address multi-scale challenges in object detection, driven by its intuitive appeal [29, 30, 31, 32, 33]. Approaches such as SNIP [34] and SNIPER [35] stand out for their selection of specific scales for each resolution during multi-scale training. However, it’s noteworthy that the utilization of image pyramids in these approaches substantially escalates inference time, presenting a trade-off between accuracy and efficiency. An alternative strategy seeks to incorporate in-network feature pyramids, offering an approximation of image pyramids with reduced computational costs. Notably, the Single-Shot Detector (SSD) [36] exemplifies this concept by leveraging multi-scale feature maps from different layers. This enables the detection of objects spanning various scales at each feature layer. Despite the computational advantages, it is essential to acknowledge that bottom-up sampling methods, like those employed in SSD, may exhibit diminished accuracy when confronted with small objects. This limitation arises from the inherent inadequacy of semantic information in low-level features, particularly impacting the precision of small object detection. Balancing the trade-offs between computational efficiency and accuracy remains a critical consideration in the ongoing evolution of multi-scale object detection methodologies within the realm of deep learning. To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [11] was introduced. FPN employs a bottom-up pathway, a top-down pathway, and lateral connections to effectively merge features of diverse resolutions and scales. Building upon FPN, the Path Aggregation Network (PANet) [37] introduced a bottom-up path to enhance information fusion between different levels. Researchers at FSAF [38] identified a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF extracts the feature representation of the region of interest (RoI) at a single feature level, overlooking any valuable information present on other feature levels of the pyramid. The SA-FPN [39], amalgamating Top-Down style FPN and Bottom-Up style FPN, combines the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrated information at all levels through common operations like element-wise max or sum, these fusion mechanisms did not consider semantic differences among feature layers. EfficientDet [40] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. YOLO-ASFF [41] employed an adaptive spatial feature fusion technique, learning weights during training to filter out inconsistencies. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [42] provided insights into this structure. Moreover, [43] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [44] proposed a parallel FPN structure for object detection with bi-directional fusion, while [45] introduced a refined FPN structure. SAFNet [46] introduced Adaptive Feature Fusion and Self-Enhanced Modules. However, some problems remain in the FPN structure. These methods all use sequential local fusion, which is an indirect interaction method and ignores the loss in the information flow process. Inspired by [14], all feature map layers present meaningful information for object detection. Therefore, designing a direct interaction and global fusion structure is of great significance for multi-scale object detection.
@@ III-A Overall Framework of SAR-Net
The overall framework of SAR-Net is illustrated in Fig. 2, where the backbone and neck components are specifically tailored for SAR images. Specifically, considering that SAR image content is predominantly background, and detected objects are directional, using feature extraction networks designed for optical images directly cannot capture the directional and positional information of objects under complex background interference. Therefore, we have designed a direction-aware feature extraction backbone network, as shown in Fig. 3. It embeds our designed direction-aware attention module, reducing background interference while capturing fine-grained information, and providing effective assistance for the subsequent global fusion of multi-scale features in SAR images. Additionally, to better address the multi-scale challenge in SAR images and avoid information dilution caused by traditional FPN local fusion, we propose a unity compensation mechanism. This mechanism comprehensively fuses deep and shallow semantic features at different scales, enabling each layer to focus on global information and providing robust prior knowledge for the detection head. The unity compensation mechanism will be detailed in the next section.
@@ III-B1 Multi-scale Alignment Module
Given the crucial premise that both shallow texture and deep semantic information of objects in SAR images are important, we choose the outputs $C2$, $C3$, $C4$, and $C5$ from the backbone for global feature alignment. We take two factors into consideration for aligning feature size: the computational complexity of global fusion and the necessity of preserving the fine-grained information extracted by the backbone. The former tends to favor smaller feature maps to reduce computational load, while the latter prefers larger feature maps to preserve detailed information. To balance these two aspects, the second smallest feature size is chosen as aligned size when four features need to be aligned while the smallest feature size is selected when aligning three features. We employ average pooling and bilinear interpolation operations to adjust the feature sizes to the aligned feature size, obtaining $\mathcal{F}_{global\_align}$. This module ensures effective aggregation of information through global alignment while preserving high-resolution features of small objects. The formula is as follows:

### v2

@@ I Introduction
Object detection plays a pivotal role in remote sensing, with extensive applications in military operations, security surveillance, disaster assessment, and environmental monitoring[1, 2]. Synthetic Aperture Radar (SAR) object detection, a key technique in remote sensing, offers notable benefits in super-resolution imaging, hidden object detection, and all-weather surveillance, particularly in identifying aircraft and ships.
In recent years, with the rapid development of deep learning, an increasing number of researchers have turned their attention to design object detection methods using Convolutional Neural Networks (CNNs), yielding notable results [3, 4, 5, 6, 7, 8, 9]. However, the unique challenges posed by SAR images, characterized by their multi-scale and directional nature, coupled with abundant background details, complicate object detection tasks. These factors particularly hinder the detection of small objects amidst background noise, making the design of effective multi-scale object detection networks for SAR images a complex task. Consequently, developing strategies to accurately determine the direction and position of aircraft or ships in SAR images, despite background interference, stands as a critical challenge in this field.
Recent advancements in SAR image object detection have prominently featured deep learning approaches. Addressing the challenges of detecting objects in varying directions, Li et al. [10] developed the Spatial Frequency Feature Fusion Network. This innovative network utilizes a Feature Pyramid Network (FPN) [11] to create spatial multi-scale features, specifically for ship objects, through a top-down hierarchical approach. Additionally, it employs a polar coordinate Fourier transform to attain rotation-invariant features in the frequency domain. In another notable work, Zhang et al. [12] introduced the Quad-Feature Pyramid Network (Quad-FPN), an architectural innovation consisting of four feature pyramid networks. This design effectively confronts issues such as background noise, multi-scale object representation, and the subtle features of small boats. Complementing these advancements, Li et al. [13] developed the Adjacent Feature Fusion (AFF) module, which strategically integrates local shallow features into adjacent layers. This selective and adaptive approach significantly enhances feature fusion, leading to improved representation of objects in SAR images.
@@
Current research in SAR image object detection primarily focuses on utilizing a FPN or its variants for feature fusion, offering a partial solution to multi-scale challenges. However, a significant limitation of these methods is their lack of global information integration at each layer. As illustrated in Fig. 1 (right side), local information flow between adjacent layers leads to a dilution of information across non-adjacent layers, resulting in information loss [14]. Recognizing the importance of global information, Bai et al. [15] proposed a Global Context-Guided Feature Balance Pyramid (GC-FBP), incorporating a global context module to bridge distant relationships between ship objects and their background. Yet, their approach of simply stacking context information onto the original feature map impedes effective interplay of information. Our method, depicted on the left side of Fig. 1, diverges from traditional algorithms that primarily depend on straightforward concatenation of adjacent feature layers. Instead, we adopt a global information fusion strategy, allowing for feature integration across various scales. This approach not only enhances the effectiveness and efficiency of the fusion process but also minimally impacts information loss during transmission.
Traditional SAR object detectors, which predominantly use horizontal boxes, typically capture features of directional objects through an attention module [16] or by transforming features into the frequency domain [10]. However, these methods tend to neglect the correlation between the object and its contextual environment. This oversight often leads to ambiguous differentiation between the object and the background, with the resultant features being significantly noisy. Furthermore, these techniques generally overlook the crucial role of positional information. In contrast, our approach consider not only the object’s features but also the surrounding context. By integrating directional and positional information and embedding channel attention, our approach offers a comprehensive solution.
We have distilled the problem of SAR image object detection into a broader challenge: addressing the issues of multi-scale and directional objects in predominantly background-dominated images. We base our considerations on three key premises: (1) the significant variation in object scales in SAR images. (2) the need to consider both shallow textural and deep semantic information of objects. (3) the directional nature of objects coupled with background interference. To address these challenges, we introduce a unity compensation mechanism, which includes three components: the Multi-scale Alignment Module (MAM), the Multi-level Fusion Module (MFM), and the Multi-feature Embedding Module (MEM). MAM initially unites multi-scale features across different spatial hierarchical levels. MFM then globally merges high-level semantic information with low-level detailed information, providing a comprehensive overview of the united features. Lastly, MEM compensates for any original feature to enrich their representational information. This multifaceted approach not only improves the detection of multi-scale objects but also significantly enhances information fusion capabilities.
In addressing the third premise, we present a direction-aware attention module equipped with bidirectional attention polymerization capabilities. This innovative module employs deformable convolution and global pooling operations to extract directional and positional information along both vertical and horizontal axes of the input features. It then generates a direction-aware attention map, which is effectively integrated with the original input features. This approach adeptly tackles the challenge of accurately discerning directional and positional information amidst intricate background interference. Our contributions can be summarized in three primary points:
$\bullet$ We introduce SAR-Net, a novel framework specifically designed for SAR image object detection. SAR-Net incorporates a unity compensation mechanism that globally unites multi-scale features of objects while compensating for the original feature. This ensures that global information is inherently integrated within features of different scales.
$\bullet$ We design a direction-aware attention module with bidirectional attention polymerization capabilities. This module extracts object features while obtaining directional and positional information, enabling the identification of object rotation poses. Such a design is particularly effective in complex SAR images, where the directional attributes of objects are often obscured by significant background interference.
$\bullet$ To validate the effectiveness, generalizability, and robustness of SAR-Net, we conduct comprehensive comparative and ablation experiments using three publicly available benchmark datasets for SAR aircraft and ship detection. The results of these experiments clearly show that our SAR-Net framework outperforms other state-of-the-art methods in all three datasets.
The remainder of this paper is organized as follows. Section II reviews the development of object detection in the field of SAR images and related literature on solving multi-scale problems. Section III details our proposed framework and modules. In Section IV, the experimental setup is described, and qualitative and quantitative results are reported. Finally, conclusions are drawn in Section V.
@@ II-A SAR Object Detection
Over the past few years, object detection in SAR images has aroused rising attention. Traditional detection algorithms, adhering to Constant False Alarm Rate (CFAR) principles [17], operate on a statistical model, dynamically setting the detection threshold based on a specified false alarm rate and the statistical distribution of background clutter. Consequently, a multitude of CFAR-based algorithms have surfaced. In a recent work [18], a novel decomposition approach was introduced to scrutinize the scattering between ships and the sea background. This decomposition technique, when integrated with CFAR, proves effective in ship detection. In a different vein, Lang et al. [19] enhanced the representation of pixels by incorporating spatial and intensity information, markedly improving the distinguishability between ships and background elements. However, the advent of high-resolution SAR images has introduced challenges such as complex ground interference and textured scenes accompanying multi-scale objects. This complexity renders the calculation of significant background clutter pixels in large-scale SAR images time-consuming. In response to the challenges of multi-scale object detection in SAR images, scholars have proposed several refined CFAR object detection algorithms. Examples include the scale sliding window object detection algorithm [20], the optimized super-pixel CFAR object detection algorithm [21], and the iterative CFAR object detection algorithm [22]. For detecting objects at larger scales, Zhai et al. [23] devised a ship detection algorithm emphasizing saliency and contextual information processing. This approach excels in focusing on sizable ships and background objects with prominent features, yielding commendable results. Despite these achievements, traditional methods still exhibit limitations in detecting small objects, necessitate manual parameter configuration, and heavily rely on manual design, presenting challenges for robustness and hindering their applicability to other SAR datasets. Consequently, there is a pressing need for more adaptive and automated approaches in SAR object detection.
Owing to the inherent capacity for autonomous feature learning in deep learning, researchers have seamlessly integrated object detection in SAR images with deep learning methodologies. Shifting focus to SAR ship detection and identification, Zhang et al. [24] proposed a joint unsupervised super-resolution and SAR object detection network. This network augments object details while suppressing noise, providing valuable guidance to the detection network. In the realm of SAR aircraft detection and recognition, Zhao et al. [25] introduced an attention module, strategically fusing refined low-level texture features with high-level semantic features. This fusion strategy enhances aircraft detection rates by leveraging complementary information. Furthermore, Zhao et al. [26] devised a multi-branch dilated convolution feature pyramid method. This innovative approach minimizes redundancy while accentuating critical aircraft features through the establishment of dense connections. In a related endeavor, Zhao et al. [27] put forth a multitask learning framework for object detection (MLDet) tailored for ship detection in SAR images. This end-to-end framework encompasses tasks such as object detection, speckle suppression, and object segmentation. The collaborative execution of these tasks contributes to an enhanced detection performance. Addressing the specific challenge of large-scale SAR images, Yang et al. [28] concentrated on a multi-channel aggregation framework. This framework holistically considers both pre-processing and post-processing in algorithm optimization, leading to an overall improvement in the effectiveness of object detection in large-scale SAR images. The comprehensive strategy employed by these studies underscores the effectiveness of deep learning paradigms in advancing the state-of-the-art in SAR object detection and recognition.
Although progress has been made, scale variation across object instances is still one of the most knotty problems in SAR object detection.
@@ II-B Multi-scale Object Detection
Within the domain of deep learning, the application of image pyramids initially emerged as a solution to address multi-scale challenges in object detection, driven by its intuitive appeal [29, 30, 31, 32, 33]. Approaches such as SNIP [34] and SNIPER [35] stand out for their selection of specific scales for each resolution during multi-scale training. However, it’s noteworthy that the utilization of image pyramids in these approaches substantially escalates inference time, presenting a trade-off between accuracy and efficiency. An alternative strategy seeks to incorporate in-network feature pyramids, offering an approximation of image pyramids with reduced computational costs. Notably, the Single-Shot Detector (SSD) [36] exemplifies this concept by leveraging multi-scale feature maps from different layers. This enables the detection of objects spanning various scales at each feature layer. Despite the computational advantages, it is essential to acknowledge that bottom-up sampling methods, like those employed in SSD, may exhibit diminished accuracy when confronted with small objects. This limitation arises from the inherent inadequacy of semantic information in low-level features, particularly impacting the precision of small object detection. Balancing the trade-offs between computational efficiency and accuracy remains a critical consideration in the ongoing evolution of multi-scale object detection methodologies within the realm of deep learning.
To address the semantic limitations in low-level features, the Feature Pyramid Network (FPN) [11] was introduced. FPN employs a bottom-up pathway, a top-down pathway, and lateral connections to effectively merge features of diverse resolutions and scales. Building upon FPN, the Path Aggregation Network (PANet) [37] introduced a bottom-up path to enhance information fusion between different levels. Researchers at FSAF [38] identified a flaw in FPN’s heuristic-guided feature selection, highlighting the need for improvement. In contrast, FSAF extracts the feature representation of the region of interest (RoI) at a single feature level, overlooking any valuable information present on other feature levels of the pyramid. The SA-FPN [39], amalgamating Top-Down style FPN and Bottom-Up style FPN, combines the strengths of both, emerging as a more accurate module for perceiving scale variations. While previous methods integrated information at all levels through common operations like element-wise max or sum, these fusion mechanisms did not consider semantic differences among feature layers. EfficientDet [40] addressed this limitation by introducing a repeatable module (BiFPN) to enhance information fusion efficiency between different levels. YOLO-ASFF [41] employed an adaptive spatial feature fusion technique, learning weights during training to filter out inconsistencies. Recognizing that the feature maps in a pyramid predominantly consist of single-level features, M2Det [42] provided insights into this structure. Moreover, [43] extended FPN with the Asymptotic Feature Pyramid Network (AFPN) to facilitate interaction across non-adjacent layers. To overcome FPN’s limitations in detecting large objects, [44] proposed a parallel FPN structure for object detection with bi-directional fusion, while [45] introduced a refined FPN structure. SAFNet [46] introduced Adaptive Feature Fusion and Self-Enhanced Modules.
However, some problems remain in the FPN structure. These methods all use sequential local fusion, which is an indirect interaction method and ignores the loss in the information flow process. Inspired by [14], all feature map layers present meaningful information for object detection. Therefore, designing a direct interaction and global fusion structure is of great significance for multi-scale object detection.
@@ III-A Overall Framework of SAR-Net
The overall framework of SAR-Net is illustrated in Fig. 2, where the backbone and neck components are specifically tailored for SAR images. Specifically, considering that SAR image content is predominantly background, and detected objects are directional, using feature extraction networks designed for optical images directly cannot capture the directional and positional information of objects under complex background interference. Therefore, we have designed a direction-aware feature extraction backbone network, as shown in Fig. 3. It embeds our designed direction-aware attention module, reducing background interference while capturing fine-grained information, and providing effective assistance for the subsequent global fusion of multi-scale features in SAR images. Additionally, to better address the multi-scale challenge in SAR images and avoid information dilution caused by traditional FPN local fusion, we propose a unity compensation mechanism. This mechanism comprehensively fuses deep and shallow features at different scales, enabling each layer to focus on global information and providing robust prior knowledge for the detection head.
@@ III-B1 Multi-scale Alignment Module
Given the crucial premise that both shallow texture and deep semantic information of objects in SAR images are important, we choose the outputs $C2$, $C3$, $C4$, and $C5$ from the backbone for global feature alignment. We take two factors into consideration for aligning feature size: the computational complexity of global fusion and the necessity of preserving the fine-grained information extracted by the backbone. The former tends to favor smaller feature maps to reduce computational load, while the latter prefers larger feature maps to preserve detailed information. To balance these two aspects, the second smallest feature size is chosen as aligned size when four features need to be aligned while the smallest feature size is selected when aligning three features.
We employ average pooling and bilinear interpolation operations to adjust the feature sizes to the aligned feature size, obtaining $\mathcal{F}_{global\_align}$. This module ensures effective aggregation of information through global alignment while preserving high-resolution features of small objects. The formula is as follows: